OpenStudy (sugarapples):

evaluate each function if g(a)= a^2-2a+1,find g(3) please help

jimthompson5910 (jim_thompson5910):

g(a)= a^2-2a+1 g(3)= (3)^2-2(3)+1 ... replace every copy of 'a' with 3 g(3) = ???